ABSORPTIONS BETWEEN 3000 AND 5500 cm^{−1} OF NORMAL AND OXYGEN-18 ENRICHED O_{4}^{+} AND O_{4}^{−} TRAPPED IN SOLID NEON

\underline{MARILYN E. JACOX} and WARREN E. THOMPSON, Optical Technology Division, National Institute of Standards and Technology, Gaithersburg, MD 20899-8441.

Very recently, Ricks, Douberly, and Duncan^{a} have assigned absorptions in the 3000 to 4300 cm^{−1} spectral region to combination bands of gas-phase cyc-O_{4}^{+}. Other experiments by Kelley, Robertson, and Johnson^{b} identified vibronic bands between 4000 and 5300 cm^{−1} of gas-phase O_{4}^{−} complexed with an argon atom. Absorptions corresponding to these bands have been observed in the present experiments for both cyc-O_{4}^{+} and O_{4}^{−}, as well as for several of their isotopologues trapped at 4.3 K in solid neon. The results will be compared with the gas-phase data, and possible assignments will be considered taking into account the results of isotopic substitution.

aA. M. Ricks, G. E. Douberly, and M. A. Duncan, Int. J. Mass Spectrom., doi:10.1016/j.ijms.2009.01.009 (2008).

bJ. A. Kelley, W. H. Robertson, and M. A. Johnson, Chem. Phys. Lett. 362, 255 (2002).